OpenStudy (marmar10):

if p(t)= s'(t), p(t)= e^5-t^2 and s(0)=0 find s(1)

OpenStudy (anonymous):

\[ s(1)-s(0) = \int_0^1s'(t)~dt \]

OpenStudy (jhannybean):

FToC*

OpenStudy (marmar10):

what is the first step? im confused by these problems :/

OpenStudy (jhannybean):

Why dont you integrate first.

OpenStudy (anonymous):

Substitute...\[ s(1) - 0 = \int_0^1e^5-t^2~dt \]

OpenStudy (anonymous):

\[ s(1) = e^5t-\frac{t^3}3\Bigg|_0^1 \]

OpenStudy (marmar10):

is the answer just \[e ^{5}- \frac{ 1 }{ 3 }\]

OpenStudy (noelgreco):

Yes

OpenStudy (marmar10):

ok thank you everyone :)